OpenStudy (anonymous):

Simplify the expression. (-1+6i) (-4+2i)

OpenStudy (lgbasallote):

FOIL it \[(-1 \times -4) + (6i \times -4) + (-1 \times 2i) + (6i \times 2i)\]

OpenStudy (lgbasallote):

what do you get from that?

OpenStudy (anonymous):

(-8-26i)

OpenStudy (anonymous):

-5+8i